If the hypotenuse of a right triangle is 30 cm long and one leg is twice as long as the other, how long are the legs of the triangle?
Answer by checkley77(12844)

a^2+b^2=c^2
a^2+(2a)^2=30^2
a^2+4a^2=900
5a^2=900
a^2=900/5
a^2=180
a=sqrt180
a=13.4164 ans. for the short leg.
2*13.4164=26.8328 ans. for the long leg.
Proof
13.4164^2+26.8328^2=900
180+720=900
900=900